Fermi National Laboratory
 

Accelerator Update for July 11, 2007 - July 13, 2007

 

Wednesday July 11

The day shift began with most of the accelerators and all the experiments in access while the Collider Detector at Fermilab (CDF) made some necessary repairs to their detector. Only the Recycler remained running, which kept the Main Injector (MI) locked up. Operations monitored the stash (338.5E10).

Linac technicians began work on a RF station (LRF4) at 8:28 AM.

Cryo system experts began work on the Antiproton Sourc (Pbar) Debuncher cryo system at 8:50 AM.

All Linac work ended and Operations began sending beam to the Linac absorber at 1:30 PM.

Operations reported at 2:55 PM that Preaccelerator experts switched to the I- Source.

Kicker experts reported at 3:03 PM that they had replaced the trigger units and filter capacitors on the MI-10 KPS4A & B.

Linac experts began a LRF5 study at 3:09 PM.

At 3:16 PM, Operations reported that 4E10 of the stashed antiprotons were lost and then the Recycler suffered an emittance blow up. Experts investigated.

Operations began the Pbar turn on at 3:55 PM.

A new Meson MTest experiment (T958), which will run parasitically, received its approval to take beam at 4:52 PM.

Operations turned on MI at 5:25 PM.

Operations turned on NuMI and MiniBooNE at 5:29 PM.

Operations reported at 6 PM that they were having a problem with a MI kicker (KPS1B). Operations called in an expert. Operations began sending beam to the Booster absorber.

Operations established beam to MiniBooNE at 6:30 PM.

Operations established beam to NuMI at 8:06 PM.

Kicker experts reported at 9:24 PM that they had completed their work on KPS1B. They authorized Operations to lower the reservoir voltage if it tripped again.

Pbar experts began a reverse proton study at 10:05 PM.

At 10:06 PM, KPS1B tripped off. It reset.

CDF came out of access at 10:35 PM.

Operations put the TeV through a dry squeeze at 10:44 PM.

Cryo system experts reported at 10:54 PM that the Pbar cryo system would be ready in about five hours.

Operations put the TeV through a wet squeeze at 11:42 PM.

Thursday July 12

The midnight shift began with the TeV in a wet squeeze, with Operations monitoring the stash (317.9E10), with Pbar conducting reverse proton studies, and with NuMI and MiniBooNE taking beam.

The wet squeeze ended at 12:55 AM.

The reverse proton studies ended and Pbar began stacking at 1:11 AM.

At 1:44 AM, Operations reported that MIRF15 tripped off. Operators couldn’t discover why it had tripped off, nor could they recover the station. Operators decided to bypass it and try to turn it on later.

At 2:22 AM, Operations reported that 35E10 of antiprotons were lost from the Recycler stash. The emittance grew very large. Operators called in an expert.

At 3:17 AM, the 750 KeV line buncher tripped off. It reset.

Operations put the TeV through a dry squeeze at 4:10 AM.

Operations successfully turned on MIRF15 at 4:47 AM.

Operations put the TeV into shot setup at 6:03 AM.

Operations established store 5549 at 8:40 AM with an initial luminosity of 139.5E30.

At 9:25 AM, Operations reported a problem with the Meson MTest safety system. Interlock experts investigated.

Operations established beam to MTest at 10:28 AM.

Interlock experts reported at 10:29 AM that they found and replaced a bad power supply for the Meson MTest safety system.

Experts began a MI/MiniBooNE study at 1:08 PM.

At 2:47 PM, the Linac low-energy buncher tripped off. It reset.

Operations reported at 4:06 PM that the Recycler had stashed antiprotons three times during the day shift.

At 4:34 PM, all the devices in the Pbar AP-1 line tripped off. Simultaneously, there was an indication that feeder 89 had opened. Operations contacted the Duty Electrician.

At 4:57 PM, LRF2 tripped off. An expert had it back up and running within 15 minutes.

The Run Coordinator reported at 5 PM that feeder 89-1 (F sector) tripped off on a ground fault.

The Run Coordinator reported at 5:15 PM that he and the Duty Electrician would hipot the F sector portion of feeder 89 and try to find where the problem was, which would take about two hours.

Operations disabled MTest for the night at 5:19 PM. They had been in access for most of the afternoon due to the feeder 89 problem.

The Run Coordinator reported at 8:55 PM that a P3 line dipole magnet power supply (HP3US) was the cause of the feeder 89 problem. Experts disconnected the power supply.

Pbar resumed stacking at 9:15 PM.

Operations reported at 11:58 PM that the Recycler had stashed once during the evening shift.

Friday July 13

The midnight shift began with Operations monitoring store 5549 (38.2E30) and the stash (159.8E10), with Pbar stacking (50.8E10), and with NuMI taking beam.

At 12:26 AM, BRF16 tripped off. It reset.

At 12:31 AM, a Booster vacuum pump tripped off and took down BRF7 & 8. Everything reset.

At 3:08 AM, the NuMI beam permit tripped off due to beam in the kicker gap. It reset.

At 4:59 AM, the Pbar sweeper magnet tripped off. It reset.

Operations terminated store 5549 at 7:06 AM.

Operations prepared D0 for an access at 7:25 AM.

Operations reported at 8:12 AM that the Recycler had stashed four times during the midnight shift.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/13/2007    email Fermilab